# Breaking degeneracies in exoplanetary parameters through self-consistent atmosphere-interior modelling

### v2

@@ 6 Case study : WASP-39 b, a transiting hot-Saturn
WASP-39 b is an inflated Saturn-like planet orbiting a G-type star with a period of P (Faedi et al. 2011). Radial velocity and transit photometry measurements as described by Faedi et al. (2011) indicate a mass of MJ and a radius of MJ, hence suggesting an inflated planet. Ahrer et al. (2023) detected the presence of from its strong 4.4 band from JWST NIRSpec-G395H observations, as part of the Early Release Science Program (ERS). This detection suggests a relatively high metallicity (solar metallicity). We propose to use this JWST transit spectrum to exemplify the model. Two approaches are adopted: on the one hand, we assume no spectrum and uniquely mass and radius measurements. Hence equation 14 is reduced down to the second term. On the other hand, we use the JWST ERS spectrum with the radial velocity mass measurement. In this case, equation 14 includes the first term. The objective is to understand the impact that spectral information has on constraining parameters such as the intrinsic temperature derived for example in Figure 4. We do not include a radius measurement as it is given within the transit depth at the JWST NIRSpec-G395H wavelengths. Radius measurements from for instance Faedi et al. (2011) correspond to the radius in the visible and near-infrared (R and Z band).
@@
Figure 8 shows the posterior distributions for the two scenarios. The retrieved values are given in table 4. We see that without the spectrum, the posterior distributions show a highly degenerated ensemble of potential parameters. Most notably, there are no constraints on the metallicity or the core mass. Possible values for the intrinsic temperature considering a Gaussian posterior with 1 variation are between 150K and 320K. For the case with the spectrum, we ran the retrieval based on our grid of models using a chemistry at equilibrium. Stronger constraints are obtained for the metallicity and core mass. A relatively high metallicity of 10 solar is necessary to fit adequately the band as shown in the red 100 best fits of Figure 9. A high Tint of 297.39K is required to reproduce the inflated radius, given the equally high metallicity, but also to convert CH4 into CO and CO2 leading to the observed CO2 spectral signature. The core proposed is relatively large, between 20 to 30 times earth’s mass, hence representing up to 30 of the total planetary mass. To match the higher atmospheric metallicity, the helium mass fraction in the interior is approximately 0.50. As such, considering a Solar-like helium mass fraction of 0.246 (Lodders 2019) the mass of heavy elements (core mass + excess helium) is approximately 45 of the planet’s total mass. Bloot et al. (2023) finds a core mass of 0.88 times Earth’s mass using a homogeneous model and 0.90 times Earth’s mass using a non-homogeneous core model. In both cases, JWST data was not used. Furthermore, Thorngren & Fortney (2019) use a bulk metallicity of 22 the planet’s mass to derive an atmospheric metallicity of 40.51 times solar, again without using JWST data. We now know the atmospheric metallicity to be lower and within the framework of their approach this could change how well mixed the planet is. In all cases, we find a core mass that is substantially higher with the JWST data. Without the JWST data, the result is very degenerate and would include both possibilities. Figure 9 shows the fit over-plotted on to the NIRSpec-PRISM spectrum from Rustamkulov et al. (2023). The signature is shown to be well fitted by the model. Discrepancies arise at shorter wavelengths due to the absence of clouds in the current model. The absence of photo-chemistry and in the model explains the missing absorption band around 4 m. This however has no impact on the overall fit. We however are able to reproduce the sodium signature around 0.5 m. We conclude that adding spectroscopic measurements enables us to break the degeneracies between metallicity, intrinsic temperature and core mass. The fact that the metallicity and core values found for the spectrum-constrained case are outside the 68 region of the unconstrained posterior distribution, indicates that the spectrum provides essential information that is not included with mass and radius alone. The metallicity derived here using the transit spectrum is in agreement with the 10 solar value found by the extensive range of models applied to this planet and spectrum (Powell et al. 2024; Constantinou & Madhusudhan 2024; Carone et al. 2023; Feinstein et al. 2023; JWST-ERS et al. 2023; Alderson et al. 2023). Using the same spectrum and the retrieval code TauREx from Al-Refaie et al. (2021) and used for example in Panek et al. (2023), we find a metallicity ranging from 11.99 to 12.85 times solar. This result is also in agreement with the derived atmospheric metallicity found using the developed model. Alderson et al. (2023) finds a wide range of intrinsic temperatures using three atmosphere models : Atmo (K), PHOENIX (K) and PICASO+Virga (K). This could be due to the lack of coupling between an atmosphere and interior model as is done here.
Parameter | Mass-Radius retrieval | Mass-Spectra retrieval |
@@
7 Case study : 51 Eridani b, a remote young massif-Jupiter
@@
51 Eridani b (hereafter 51 Eri b) is a massive-Jupiter planet orbiting the young ( Myr) F0 type star 51 Eridani. The high uncertainty on the age makes the interpretation of the photometric data to derive the planet’s physical parameters difficult. Identified in 2015 in high contrast imaging, on the basis of a absorption in its near infrared spectrum (Macintosh et al. 2015), using the Gemini Planet Imager (GPI, Macintosh et al. 2014). In a subsequent paper, Rajan et al. (2017) used the same GPI Integral Field Spectrograph (IFS) data to show that 51 Eri b was consistent with an effective temperature ranging between 605 and 737 K, a solar metallicity, and a surface gravity of log(g) = 3.5–4.0. Samland et al. (2017) shows a super solar metallicity at [Fe/H] = 1.03dex, a 759.51 effective temperature and a surface gravity of log(g) = 4.26. More recently, Brown-Sevilla et al. (2023) used photometric data obtained with the Spectro-Polarimetic High contrast imager for Exoplanets REsearch (SPHERE, Beuzit et al. 2019) at the VLT to constrain its gravity to log(g)= and its radius to . Unlike the previous case study described in Sect. 6, only spectrophotometric data are available. We can then reduce equation 14 down to the first term. This equally means that we have fewer observational constraints to infer the structure of the exoplanet.
@@
Using the same observation as in Brown-Sevilla et al. (2023) and Rajan et al. (2017), we extracted a YJH spectrum using SPHERE/IFS (Claudi et al. 2008) and the H spectrum using two GPI/IFS epochs. The observing conditions of the observations can be found in Annex Table 3. The extraction of a spectrum can be complicated and the uncertainties associated with this process must be kept in mind when subsequently running retrievals.
@@
The pre-processing steps consist of corrections to the sky, the flat and the background, bad pixel interpolation and wavelength calibration. These steps use the standard pipeline installed on the High Contrast Data Center, formerly known as the Sphere Data Center (Delorme et al. 2017). Regarding GPI, every raw frame was dark subtracted, flat-fielded, cleaned of correlated detector noise, bad pixel corrected, flexure corrected using the standard GPI Data Reduction Pipeline (DRP; version 1.6) documented in Perrin et al. (2014, 2016). The resulting set of frames was combined into a spectral datacube using Pyklip (Wang et al. 2015), that allowed for a more accurate wavelength and PSF recalibration in addition to frame centering.
As a post processing algorithm, we consistently use PACO ASDI
@@
(Flasseur et al. 2020a), an ADI based algorithm. It models the noise at a local scale in small patches, for all the datasets. PACO relies on the learning of local temporal and spectral correlations of the noise to provide reliable and statistically grounded signal to noise ratio map and photometry confidence interval. A full detailed explanation of the algorithm can be found in (Flasseur et al. 2018, 2020a, 2020b) and its implementation on the Data Center and spectral optimization in Chomez et al. (2023).
@@
After measuring with PACO the contrast spectrum of the planet, that is the ratio of the flux coming from the planet to the stellar PSF flux, we obtained a physical spectrum by multiplying such contrast by a synthetic spectrum of the star based on BT-NextGen AGSS09 atmospheric models (Allard et al. 2012). The stellar spectrum is normalized by 2MASS (Skrutskie et al. 2006) H magnitude. In addition to the H-band uncertainty, the final spectrum error budget also accounts for systematic uncertainties on the star PSF intensity (Wang et al. 2014), as well as for random uncertainties related to the PSF variations over the exposure and to PACO’s fit. The two GPI spectra were finally averaged to get the H-band spectrum used in the present Section. A similar procedure was applied to SPHERE data. Errobars for both SPHERE and GPI are displayed at 1.
@@
Annexe 13 shows the corner plot of the spectral retrievals using MCMC, the best fit is shown in Figure 10 and the retrieved values given in table 5. While the fit shown shows similar spectral features and close overall flux amplitude to the observed spectrum, there are notable differences below . This could be due to the presence of clouds which are not included in this current version of the model. Charnay et al. (2018) shows how clouds have an effect on emission spectra and most notably in the Y, J, H bands. Annexe 13 indicates a multinomial distribution with 3 mass solutions found with one of the mass solutions further degenerate in intrinsic temperature and metallalicity. All solutions are given in Table 5. In order to break the degeneracy, we include in the corner plot, the age estimations found by solving equation 8. The three proposed ages are to be compared to the age estimate of the star of proposed by Macintosh et al. (2015). We identify the higher mass case () as being the most compatible configuration. This configuration has a corresponding age of Myr and is hence in strong agreement with the age of the system. It is however important to mention that the planet and the star need not be coeval. However, Mordasini et al. (2017) shows that a 5 MJ would not take longer than 10 Myrs to finish it’s mass accretion with observations indicating less than 5 Myr. As such when comparing the next closest age solutions, (solutions 2 and 3) with the best match (solution 4), we see that the age of 19.5 Myr is a better match than 7 to 8 Myr.
@@
As discussed in section 3 the initial entropies used to evaluate the age are under the hot start assumption. The MCMC posterior distribution shows a core mass of 31.86ME for the best matching age case. This result, within the models approximations, indicates the presence of a core and suggests a formation with core accretion processes combined with the hot-start scenario. Such scenarios have been studied for example by Berardo & Cumming (2017); Mordasini et al. (2017). Mordasini et al. (2017) shows that it is possible for planets to undergo core accretion and have similar luminosities to those expected in a hot start or a warm start scenario. We explored the cold-start hypothesis for initial entropies, but all solutions found were below the estimated system age.
@@
The radius of 1.25RJ found here is inconsistent with the ”New nominal” radius found by Brown-Sevilla et al. (2023) of . The radius is a parameter deduced from the input parameters as defined by equation 12. The model flux is scaled according to equation 13. As such the radius deduced through the model is directly linked to the flux scaling as well as the input parameters. This is not the case with purely atmospheric retrievals where the radius becomes a free input parameter.
@@ 8 Discussion
The two examples show the advantages of a coherent model such as the one we propose here. The internal links between the mass, radius, metallicity, core size and intrinsic temperature allow for physically coherent results to be derived. Trends in parameters such as intrinsic temperature derived from mass, radius and effective temperature, need to be treated with upmost caution. The case of WASP-39b exemplifies this, although being a planet with a mass below those studied in Figure 4, it represents an interesting case where the metallicity and chemical composition derived from the spectrum drives the model towards a higher intrinsic temperature and helps explain the apparent inflated radius. WASP-39 b showed that deriving the intrinsic temperature as well as the internal structure from the mass and radius can lead to wide range of possible chemical and thermal structures. This, repeated over the range of observed exoplanets, can lead to increased uncertainties on overall trends. Using spectral data, it is possible to constrain the thermal and structural properties in a way that should pave the way for a reevaluation of planetary statistics. Thorngren et al. (2019) uses metallicity trends from Thorngren et al. (2016) which provides a constraint that corresponds to having a spectrum. The uncertainty related to this metallicity trend is however large. This leads us to believe that it is best to derive the metallicity from the spectrum on a planet-wise basis. This will become possible with the range of planets observed with JWST (Greene et al. 2019; Beichman et al. 2019) and subsequently the dedicated Ariel mission (Tinetti et al. 2020).
@@
The fits obtained for directly imaged planets need to rely on physical models that are self-consistent in order to correctly determine physical characteristics. Commonly used methods of decoupling radius from the model will potentially lead to erroneous conclusions due to non physical radii helping to adjust the overall amplitude of the spectrum. The model proposed here has the potential to be able to derive physically coherent most-likely parameters for directly imaged planets independently of their age.
@@
Three essential elements are missing from the current model, C/O ratio, H2O EOS, clouds and diffused cores. The C/O ratio is famously hard to derive; it would be interesting to evaluate the capacity to constrain this parameter within a physically coherent model. The H2O EOS is essential for exploring higher metallicity planets especially Neptune and sub-Neptune like planets. It might also have an impact on the distribution of heavy elements in exoplanet interiors should for instance metals be over-represented by the core mass. Clouds represent an important feature within the spectrum, changing the flux intensity between shorter and longer wavelengths and changing the thermal evolution of a planet. As we have shown, not including this parameter with a self-consistent model such as the one presented here, can lead to physically incoherent results in other parameters. While it may be possible to account for differences in a parametric way, a self-coherent model would be in the spirit of this current work. Finally, incorporating diffused cores and examining constraints on the extent of the core in the context of an increasing amount of exoplanet data may provide better insights into this process. As pointed out in the introduction, diffused cores provide better explanations for the Juno data than non diffused cores. While the impact on the radius is limited for older and lower intrinsic temperature planets, warmer younger planets might provide greater insight into the interior design. This addition to the model might also help better understand the distribution of heavier metals found within the core and the envelope leading to a better understanding of the core mass.

### v3

@@ 6 Case study: WASP-39 b, a transiting hot-Saturn
WASP-39 b is an inflated Saturn-like planet orbiting a G-type star with a period of P (Faedi et al. 2011). Radial velocity and transit photometry measurements as described by Faedi et al. (2011) indicate a mass of MJ and a radius of MJ, suggesting an inflated planet. Ahrer et al. (2023) detected the presence of from its strong 4.4 band from JWST NIRSpec-G395H observations, as part of the Early Release Science (ERS) program. This detection suggests a relatively high metallicity (solar metallicity). We propose to use this JWST transit spectrum in an exemplary application of the model. Two approaches are adopted: in one instance, we assume no spectrum and uniquely mass and radius measurements, and hence equation 14 is reduced to the second term; while in another, we use the JWST ERS spectrum with the radial-velocity mass measurement. In this case, equation 14 includes the first term. The objective is to understand the impact that spectral information has on the constraint of parameters such as the intrinsic temperature derived for example in Figure 4. We do not include a radius measurement as it is given within the transit depth at the JWST NIRSpec-G395H wavelengths. Radius measurements from, for example, Faedi et al. (2011) correspond to the radius in the visible and near-infrared (R and Z bands).
@@
Figure 8 shows the posterior distributions for the two scenarios. The retrieved values are given in Table 4. We see that without the spectrum, the posterior distributions show a highly degenerated ensemble of potential parameters. Most notably, there are no constraints on the metallicity or the core mass. Possible values for the intrinsic temperature considering a Gaussian posterior with 1 variation are between 150K and 320K. For the case with the spectrum, we ran the retrieval based on our grid of models using a chemistry at equilibrium. Stronger constraints are obtained for the metallicity and core mass. A relatively high metallicity of about ten times solar is necessary to adequately fit the band as shown in the best fit of Figure 9. A high Tint of 297.39K is required to reproduce the inflated radius, given the equally high metallicity, but also to convert CH4 into CO and CO2 leading to the observed CO2 spectral signature. The core proposed is relatively large, namely of between 20 and 30 times Earth’s mass, and therefore represents up to 30 of the total planetary mass. To match the higher atmospheric metallicity, the helium mass fraction in the interior is approximately 0.50. As such, considering a solar-like helium mass fraction of 0.246 (Lodders 2019), the mass of heavy elements (core mass + excess helium) is approximately 45 of the total mass of the planet. Bloot et al. (2023) finds a core mass of 0.88 times Earth’s mass using a homogeneous model and 0.90 times Earth’s mass using a non-homogeneous core model. JWST data were not used in either case. Furthermore, Thorngren & Fortney (2019) use a bulk metallicity of 22 of the planet’s mass to derive an atmospheric metallicity of 40.51 times solar, again without using JWST data. We now know the atmospheric metallicity to be lower, and within the framework of their approach this could change how well mixed the planet is. In all cases, we find a core mass that is substantially higher with the JWST data. Without the JWST data, the result is very degenerate and would include both possibilities. Figure 9 shows the fit over-plotted on the NIRSpec-PRISM spectrum from Rustamkulov et al. (2023). The signature is shown to be well fitted by the model. Discrepancies arise at shorter wavelengths due to the absence of clouds in the current model. The absence of photo-chemistry and in the model explains the missing absorption band around 4 m. Nevertheless, this has no impact on the overall fit. However, we are able to reproduce the sodium signature at around 0.5 m. We conclude that adding spectroscopic measurements enables us to break the degeneracies between metallicity, intrinsic temperature, and core mass. The fact that the metallicity and core values found for the spectrum-constrained case are outside the 68 region of the unconstrained posterior distribution indicates that the spectrum provides essential information that is not included with mass and radius alone. The metallicity derived here using the transit spectrum is in agreement with the approximately ten times solar value found by the extensive range of models applied to this planet and spectrum (Powell et al. 2024; Constantinou & Madhusudhan 2024; Carone et al. 2023; Feinstein et al. 2023; JWST-ERS et al. 2023; Alderson et al. 2023). Using the same spectrum and the retrieval code TauREx from Al-Refaie et al. (2021) —and used for example in Panek et al. (2023)—, we find a metallicity ranging from 11.99 to 12.85 times solar. This result is also in agreement with the derived atmospheric metallicity found using the developed model. Alderson et al. (2023) find a wide range of intrinsic temperatures using three atmosphere models: Atmo (K), PHOENIX (K), and PICASO+Virga (K). This could be due to the lack of coupling between the atmosphere model and the interior model, to which particular attention was paid in the present study.
Parameter | Mass–radius retrieval | Mass–spectra retrieval |
@@
7 Case study: 51 Eridani b, a young remote massive-Jupiter
@@
51 Eridani b (hereafter 51 Eri b) is a massive-Jupiter planet orbiting the young ( Myr) F0 type star 51 Eridani. The high uncertainty on its age makes it difficult to interpret the photometric data in order to derive the planet’s physical parameters. 51 Eridani b was dentified in 2015 in high-contrast imaging on the basis of a absorption in its near-infrared spectrum (Macintosh et al. 2015) using the Gemini Planet Imager (GPI, Macintosh et al. 2014). In a subsequent paper, Rajan et al. (2017) used the same GPI Integral Field Spectrograph (IFS) data for 51 Eri b, finding an effective temperature ranging between 605 and 737 K, solar metallicity, and a surface gravity of log(g) = 3.5–4.0. Samland et al. (2017) show a supersolar metallicity at [Fe/H] = 1.03dex, a 759.51 effective temperature, and a surface gravity of log(g) = 4.26. More recently, Brown-Sevilla et al. (2023) used photometric data obtained with the Spectro-Polarimetic High-contrast imager for Exoplanets REsearch (SPHERE, Beuzit et al. 2019) at the VLT to constrain its gravity to log(g)= and its radius to . Unlike the previous case study described in Sect. 6, only spectrophotometric data are available. We can then reduce equation 14 down to the first term. This equally means that we have fewer observational constraints to infer the structure of the exoplanet.
@@
Using the same observation as in Brown-Sevilla et al. (2023) and Rajan et al. (2017), we extracted a YJH spectrum using SPHERE/IFS (Claudi et al. 2008) and the H spectrum using two GPI/IFS epochs. The observing conditions of the observations can be found in Table 5 of the Appendix. The extraction of a spectrum can be complicated and the uncertainties associated with this process must be kept in mind when subsequently running retrievals.
@@
The pre-processing steps consist of corrections to the sky, the flat field, and the background, as well as bad pixel interpolation and wavelength calibration. These steps use the standard pipeline of the High Contrast Data Center, formerly known as the Sphere Data Center (Delorme et al. 2017). Regarding GPI, every raw frame was dark subtracted, flat-fielded, cleaned of correlated detector noise, bad pixel corrected, and flexure corrected using the standard GPI Data Reduction Pipeline (DRP; version 1.6) documented in Perrin et al. (2014, 2016). The resulting set of frames was combined into a spectral datacube using Pyklip (Wang et al. 2015), which allowed a more accurate wavelength and PSF recalibration in addition to frame centring.
As a post-processing algorithm, we consistently use PACO ASDI
(Flasseur et al. 2020a)
for all datasets. PACO ASDI
@@
is an ADI-based algorithm that models the noise at a local scale in small patches. It relies on the learning of local temporal and spectral correlations of the noise to provide a reliable and statistically grounded signal-to-noise-ratio map and photometry confidence intervals. A full and detailed explanation of the algorithm can be found in Flasseur et al. (2018, 2020a, 2020b) and its implementation on the Data Center and spectral optimization in Chomez et al. (2023).
@@
After measuring the contrast spectrum of the planet with PACO, that is the ratio of the flux coming from the planet to the stellar PSF flux, we obtained a physical spectrum by multiplying this contrast by a synthetic spectrum of the star based on BT-NextGen AGSS09 atmospheric models (Allard et al. 2012). The stellar spectrum is normalised by 2MASS (Skrutskie et al. 2006) H magnitude. In addition to the H-band uncertainty, the final spectrum error budget also accounts for systematic uncertainties on the star PSF intensity (Wang et al. 2014), as well as for random uncertainties related to the PSF variations over the exposure and to PACO’s fit. The two GPI spectra were finally averaged to get the H-band spectrum used in the present section. A similar procedure was applied to SPHERE data. Error bars for both SPHERE and GPI are displayed at 1.
@@
Appendix 13 shows the corner plot of the spectral retrievals using MCMC, the best fit is shown in Figure 10, and the retrieved values are given in Table 3. While this fit shows similar spectral features and an overall flux amplitude that is close to the observed spectrum, there are notable differences below . These could be due to the presence of clouds, which are not included in the current version of the model. Charnay et al. (2018) show how clouds have an effect on emission spectra and most notably in the Y, J, and H bands. Appendix 13 indicates a multi-nomial distribution with three mass solutions found, one of which is further degenerate in intrinsic temperature and metallicity. All solutions are given in Table 3. In order to break the degeneracy, we include the age estimations found by solving equation 8 in the corner plot. The three proposed ages are to be compared to the age estimate of the star of proposed by Macintosh et al. (2015). We identify the higher-mass case () as being the most compatible configuration. This configuration has a corresponding age of Myr and is hence in strong agreement with the age of the system. It is however important to mention that the planet and the star need not be coeval. However, Mordasini et al. (2017) show that a 5 MJ would take no longer than 10 Myr to finish its mass accretion with observations indicating less than 5 Myr. As such, when comparing the next closest age solutions (solutions 2 and 3) with the best match (solution 4), we see that the age of 19.5 Myr is a better match than 7 to 8 Myr.
@@
As discussed in section 3, the initial entropies used to evaluate the age are under the assumption of a hot start. The MCMC posterior distribution shows a core mass of 31.86ME for the best matching age case. This result, within the models approximations, indicates the presence of a core and suggests a formation with core-accretion processes combined with the hot-start scenario. Such scenarios have been studied for example by Berardo & Cumming (2017) and Mordasini et al. (2017). Mordasini et al. (2017) show that it is possible for planets to undergo core accretion and have similar luminosities to those expected in both hot-start and warm-start scenarios. We explored the cold-start hypothesis for initial entropies, but all solutions found were below the estimated system age.
@@
The radius of 1.25RJ found here is inconsistent with the ‘New nominal’ radius found by Brown-Sevilla et al. (2023) of . The radius is a parameter deduced from the input parameters as defined by equation 12. The model flux is scaled according to equation 13. As such, the radius deduced through the model is directly linked to the flux scaling as well as the input parameters. This is not the case with purely atmospheric retrievals, where the radius becomes a free input parameter.
@@ 8 Discussion
The two examples show the advantages of a coherent model such as the one proposed here. The internal links between the mass, radius, metallicity, core size, and intrinsic temperature allow physically coherent results to be derived. Trends in parameters, such as intrinsic temperature derived from mass, radius, and effective temperature, need to be treated with utmost caution. The case of WASP-39b exemplifies this; albeit a planet with a mass below those studied in Figure 4, it represents an interesting case where the metallicity and chemical composition derived from the spectrum drives the model towards a higher intrinsic temperature and helps explain the apparent inflated radius. WASP-39 b showed that deriving the intrinsic temperature as well as the internal structure from the mass and radius can lead to a wide range of possible chemical and thermal structures. Repeating this over the range of observed exoplanets can lead to increased uncertainties on overall trends. Using spectral data, it is possible to constrain the thermal and structural properties in a way that should pave the way for a re-evaluation of planetary statistics. Thorngren et al. (2019) use metallicity trends from Thorngren et al. (2016), which provides a constraint that corresponds to having a spectrum. However, the uncertainty related to this metallicity trend is large. This leads us to believe that it is best to derive the metallicity from the spectrum on a planet-wise basis. This will become possible with the range of planets observed with JWST (Greene et al. 2019; Beichman et al. 2019) and subsequently the dedicated Ariel mission (Tinetti et al. 2020).
@@
The fits obtained for directly imaged planets need to rely on physical models that are self-consistent in order to correctly determine physical characteristics. Commonly used methods for decoupling the radius from the model will potentially lead to erroneous conclusions due to the contribution of nonphysical radii in adjusting the overall amplitude of the spectrum. The model proposed here has the potential to provide the most physically coherent parameters for directly imaged planets, even without prior knowledge of their age.
@@
Four essential elements are missing from the current model, namely the C/O ratio, the H2O EOS, clouds, and a diffuse core. The C/O ratio is famously hard to derive; it would be interesting to evaluate our capacity to constrain this parameter within a physically coherent model. The H2O EOS is essential for exploring higher-metallicity planets, and especially Neptune and sub-Neptune-like planets, and might also have an impact on the distribution of heavy elements in exoplanet interiors should for instance metals be over-represented by the core mass. Clouds represent an important feature within the spectrum, changing the flux intensity between shorter and longer wavelengths and changing the thermal evolution of a planet. As we have shown, not including this parameter with a self-consistent model such as the one presented here can lead to physically incoherent results for other parameters. While it may be possible to account for differences in a parametric way, a self-coherent model would be in the spirit of the current work. Finally, incorporating diffused cores and examining constraints on the extent of the core in the context of an increasing amount of exoplanet data may provide better insights into this process. As pointed out in Section 1, diffuse cores provide better explanations for the Juno data than non-diffuse cores. While the impact on the radius is limited for older planets and those with lower intrinsic temperatures, warmer, younger planets might provide greater insight into the interior design. This addition to the model might also help us to better understand the distribution of heavier metals found within the core and the envelope, leading to a better understanding of the core mass.